eniment/Subject 05 Date Name Lab Partner Course & Section No. Locker Desk No. Part2 mass ater trot Colume water ddeliver mass water delvere Temperature 23 C CRC 9975409 empty battle 26.4389 9.qu7mL mass of botlet lo mLwat er 9.9239 36. 3619 bottle t 2omL water 9.111 u6.lu09 9-803 mL 9.803 mL bottle +3ML water 55.9199 bottle + uo mし 9.8839 66.8029 9.907 mL Wat-tv 75-5749 9.772 9 9. su69 9-796 mレ bottlet 50 mL 35.4209 bottle + 60 mL 9.370 mL Signature Date Witness/TA Date THE HAYDEN-MCNEIL STUDENT LAB NOTEBOOK Note: Insert Divider Under Copy Sheet Before Writing
Part II: Graduated Cylinder Calibration 1. Calculate the average and standard deviation for the mass of water contained in the cylinder. 2. Find the density of water at the temperature you recorded in the Handbook of Chemistry Physics. Use this density to convert the mean value of the mass of water and its standard devia tion to the corresponding volume. a. What does the standard deviation tell you about the precision of the beaker? You will use your data to support your choices in your discussion. ana
Procedure Part B: Density of Coke and Diet Coke using Calibrated Glassware 3. What is the percent error? What does the percent error tell you about the accuracy of your mea- surements? Classify the accuracy as good, fair, or poor as you did before.
Answer #1

Average (X) = (9.947 + 9.803 + 9.803 + 9.907 + 9.796 + 9.870)/6 = 9.854

Standard deviation = Σα- Χ2 X)2 η.

Here, x are the values, X is average, n is total number

x (x-X)2 [ Here, X = 9.854]
9.947 (9.947 - 9.854)2 = 0.008649
9.803 (9.803 - 9.854)2 =0.002601
9.803 (9.803 - 9.854)2 =0.002601
9.907 (9.907 - 9.854)2 =0.002809
9.796 (9.796 - 9.854)2 =0.003364
9.870 (9.870 - 9.854)2 =0.000256
(x - X) = 0.02028

Σα- X 0.02028 0.0581 6 η.

Using density = 0.9975048 g/mL

Average volume = Average mass/density = 9.854/0.9975048 = 9.879 mL

Standard deviation in volume = 0.0581/0.9975048 = 0.0528

Answer a)

As standard deviation is low, it means the the experimental method is precise as the difference between largest measured value and smallest measure value is not so high

Answer 3

Percent error = [(Experimental value - Accepted value)/Accepted value]*100

More is percent error lesser is the accuracy

Good accuracy is when accepted value and experimental values are almost same

Low accuracy is when accepted value and experimental values are very different